# A New Causal Rule Learning Approach to Interpretable Estimation of Heterogeneous Treatment Effect

### v2

@@ Appendix D Descriptive statistics of ASD variables
atsize0 (mean (SD)) 9.41 (6.39) 17.31 (8.61) $<$0.001 1.042
sysbp (mean (SD)) 203.04 (13.33) 99.58 (12.07) $<$0.001 0.272
@@
oroom 0.22 (0.42) 0.31 (0.46) $<$0.001 0.207
occnum 1.00 (0.13) 0.10 (0.30) $<$0.001 3.858
periop 1.12 (0.37) 3.09 (0.82) $<$0.001 3.111
@@
HFD1Y 360.05 (4.75) 352.49 (6.17) $<$0.001 1.373
Appendix E Propensity score matching of the ASD data
Figure 1: Detials of propensity score matching on the ASD data, including the density of propensity scores before (a) and after (b) matching and love plot (c) of both treatment groups.
We use R package MatchIt [21] to perform the propensity score matching procedure on the pre-treatment variables of the ASD data. Note that we recode the negative treatment to 00 instead of $-1$. The procedure involves three core aspects: propensity score estimation, data matching, and balance diagnostics. We measure the distance between samples in terms of propensity scores and try various combinations of propensity score estimation models and matching methods. Our final choice goes to a covariate balancing propensity score [25] model with the following implementation formula to estimate propensity scores, by which the samples are stratified into subclasses for matching:
@@ Appendix E Propensity score matching of the ASD data
This combination is able to achieve i) good overlap within each subclass, see Figure 1(b). The improvement of overlap before (Figure 1(a)) and after (Figure 1(b)) the matching is obvious; ii) balanced covariates between two treatment groups. Before the matching, there are only five variables that are considered balanced between the two groups, having a standardized mean difference (SMD) within 10%. After the matching, all covariates are balanced other than the square root of atsize0 (sq.atsize0), whose SMD value is slightly above the threshold (0.1005). Here, we consider this difference to be acceptable as it is what we can best achieve from all attempts. See the love plot in Figure 1(c) for the comparison of covariates balance between treatments before and after the matching; iii) least observations discarded. The matching process retains a total of 2554 observations, with 2180 for positive treatment (PIC) and 374 for negative treatment (MISC). We then apply CRL on this matched data set.

### v3

@@ Appendix F Descriptive statistics of ASD variables
atsize0 (mean (SD)) 9.41 (6.39) 17.31 (8.61) <<0.001 1.042
sysbp (mean (SD)) 203.04 (13.33) 99.58 (12.07) <<0.001 0.272
@@
oroom 0.22 (0.42) 0.31 (0.46) <<0.001 0.207
occnum 1.00 (0.13) 0.10 (0.30) <<0.001 3.858
periop 1.12 (0.37) 3.09 (0.82) <<0.001 3.111
@@
HFD1Y 360.05 (4.75) 352.49 (6.17) <<0.001 1.373
Appendix G Propensity score matching of the ASD data
Figure 1: Details of propensity score matching on the ASD data, including the density of propensity scores before (a) and after (b) matching and the love plot (c) of both treatment groups.
We use the R package MatchIt [ho2011package] to perform the propensity score matching procedure on the pre-treatment variables of the ASD data. Note that we recode the negative treatment to 0 instead of 1-1. The procedure involves three core aspects: propensity score estimation, data matching, and balance diagnostics. We measure the distance between samples in terms of propensity scores and try various combinations of propensity score estimation models and matching methods. Our final choice goes to a covariate balancing propensity score [imai2014covariate] model with the following implementation formula to estimate propensity scores, by which the samples are stratified into subclasses for matching:
@@ Appendix G Propensity score matching of the ASD data
This combination is able to achieve
i) good overlap within each subclass, see Figure 1(b). The improvement of overlap before (Figure 1(a)) and after (Figure 1(b)) the matching is obvious; ii) balanced covariates between the two treatment groups. Before the matching, there are only five variables that are considered balanced between the two groups, having a standardized mean difference (SMD) within 10%. After the matching, all covariates are balanced other than the square root of atsize0 (sq.atsize0), whose SMD value is slightly above the threshold (0.1005). Here, we consider this difference to be acceptable as it is what we can best achieve from all attempts. See the love plot in Figure 1(c) for the comparison of covariate balance between treatments before and after the matching; iii) the fewest observations discarded. The matching process retains a total of 2554 observations, with 2180 for positive treatment (PIC) and 374 for negative treatment (MISC). We then apply CRL on this matched data set.